OpenStudy (anonymous):

The graph of y=5x^4-x^5 has a point of inflection at?

OpenStudy (turingtest):

points of inflection occur when f''(x)=0 y=5x^4-x^5 y'=20x^3-5x^4 y''=30x^2-20x^3=10x^2(3-2x)=0 so the inflection points are: x={0,3/2}